# Electromagnetic Property Sensing in ISAC with Multiple Base Stations: Algorithm, Pilot Design, and Performance Analysis

### v1

@@ V-A Joint Design of Multi-Subcarrier Pilots
As shown in Fig. 2, the different subcarriers provide distinct pilot patterns in region $D$. However, due to relatively small number of the UE antennas in practice, the stripes of the patterns generated by the same UE are roughly in the same direction that is mainly determined by the position of the UE. Thus, a single UE may not be sufficient to generate pilots for the reconstruction of RPCD in $D$. To address this problem, we require multiple users at different positions to provide diverse views and resolutions of RPCD. Besides, since UE-2 is closer to $D$, the near-field effect is more prominent and the pilot pattern stripes are not parallel. Closer proximity leads to more diversity in the pattern style, which is more beneficial to RPCD reconstruction. Despite the fact that each UE designs its pilots independently in (43), the pilot patterns of different UEs are inherently distinct and with low mutual coherence when the UEs are in different positions.
@@ V-B Material Identification Methodology
In order to determine the material of the target, we first need to distinguish between the parts of region $D$ occupied by different materials. To accomplish this, we utilize the density-based spatial clustering of applications with noise (DBSCAN) algorithm to divide the sampling points in $D$ into several categories [46]. DBSCAN is an unsupervised algorithm with strong generalization of clusters for different shapes and sizes, and is guaranteed to converge regardless of the input data distribution. The approach to determining the DBSCAN hyper-parameters has been discussed in [47]. Since the relative permittivity and conductivity have different dimensions, we adopt the non-dimensional and scale-invariant Mahalanobis distance as the metric in the DBSCAN clustering algorithm [48]. The number of clusters is self-adaptive and represents the estimated number of material categories. The cluster centroid of the air, which represents the average permittivity and conductivity values of the air, is typically close to $(1,0)$. Conversely, the cluster centroids of the target materials, which individually represent their average permittivity and conductivity, are generally positioned significantly away from $(1,0)$.
@@ VI Simulation Results and Analysis
Suppose the prior knowledge determines that the target is located within the region $D=[-1,1]\times[-1,1]$ $\mathrm{m}^{2}$. A total of $U=16$ UEs are randomly located in a circular region of radius $10$ m centered at the origin but outside region $D$. We consider a ULA transmitter equipped with $N_{t}=8$ antennas at each UE. The power budget is the same for all UEs and is evenly allocated to all subcarriers. For each UE, $I=16$ pilot OFDM symbols are transmitted in each subcarrier. Assume $4$ BSs are located in sequence at $(100,0)$ m, $(0,100)$ m, $(-100,0)$ m, and $(0,-100)$ m, respectively. Each BS is equipped with $N_{r}=64$ ULA antennas parallel to the $x$ direction. We assume that the noise at each BS is subject to the same zero-mean circularly symmetric complex Gaussian distribution with the same covariance. The SNR is then defined as the ratio of the overall received signal power from all BSs to the overall noise power. The inter-antenna spacing for both the UEs and the BSs is set as $0.005$ m. The central frequency is $f_{c}=30$ GHz and the frequency spacing of OFDM subcarriers is $\Delta f=120$ KHz. We choose the number of sampling points in region $D$ as $M=64\times 64=4096$. Each sampling point in $D$ can be regarded as a single pixel in an image, i.e., the RPCD image consists of $64\times 64=4096$ pixels.
@@
(b) Reconstrcted relative permittivity
@@
(c) Reconstrcted relative permittivity
@@
(e) Reconstrcted conductivity with SNR = 0 dB
(f) Reconstrcted conductivity with SNR = 30 dB
Figure 3: RPCD reconstruction results versus SNR with $L=4$ BSs and $K=20$ subcarriers. Unit of conductivity is S/m.
Figure 4: NMSE of PRCD versus SNR with $K=20$.
@@ VI-A RPCD Reconstruction Performance versus SNR
To illustrate the RPCD reconstruction results vividly, we present the reconstructed images of the target based on relative permittivity and conductivity, respectively, with $L=4$, $K=20$, and SNR $=0$ dB or $30$ dB in Fig. 3. The target “THU” is the abbreviation of “Tsinghua University”, and the three letters are composed of wood, chipboard, and plasterboard, respectively. It is seen from Fig. 3 that, the reconstructed RPCD can reflect the general shape of the target. The RPCD reconstructed at SNR $=30$ dB is much more accurate to demonstrate the target’s shape compared to the RPCD reconstructed at SNR $=0$ dB. Moreover, a higher SNR value results in more accurate reconstructed values of relative permittivity and conductivity, which leads to better representation of the target’s real EM property.
We explore the NMSE of reconstructed RPCD versus the SNR, as shown in Fig. 4. We set the number of subcarriers to be $K=20$ and sequentially select several BSs and fuse their data at the feature level to enhance the overall accuracy of the reconstructed RPCD. It is seen from Fig. 4 that, the NMSE decreases with the increase of SNR in all cases, and the best performance is achieved when $L=4$. The NMSE decreases fast in the beginning when the SNR increases from 00 dB to $10$ dB. When the SNR reaches a threshold of approximately $20$ dB, the NMSE reduces to an error floor and remains largely unchanged. At this level, the limitation on RPCD reconstruction quality is no longer due to noise at the BS but rather to the insufficient information about the target captured in the general sensing matrix. As signals from more BSs are utilized, more information of the target is integrated, which leads to a smaller error floor in RPCD reconstruction.
@@ VI-B RPCD Reconstruction Performance versus Bandwidth
We investigate the NMSE of reconstructed RPCD versus the bandwidth by changing the number of subcarriers $K$ while keeping the frequency step $\Delta f$ unchanged, as shown in Fig. 6. It is seen that, NMSE decreases with the increase of $K$ in all cases, and the best performance is achieved when the signals from all 4 BSs are used. This is because the increased bandwidth enhances the diversity of the pilot patterns and increases the number of sensing equations. Consequently, more information about RPCD is extracted, leading to better reconstruction accuracy.
@@ VI-C Material Classification Accuracy versus SNR
Figure 8: Material classification accuracy versus SNR with $K=20$.
@@
To provide an example of classifying the materials in region $D$, we show the clustering results using DBSCAN in Fig. 7 with $K=20$, SNR = $30$ dB, and $L=4$. In Fig. 7, each data point represents a sampling point in region $D$, and the colors of the data points represent the classification results. It is observed that the cluster of the air is close to $(1,0)$, whereas the clusters of the target materials are notably away from $(1,0)$. Therefore, this separation enables clear differentiation between sampling points associated with air and those associated with the target materials.
Using the sampling points occupied by the target after removing the outlier points and the points associated with air, we show the material classification accuracy versus SNR with different $L$ when $K=20$ in Fig. 8. It is seen that, the classification accuracy increases with the increase of SNR for all $L$, and the best performance is achieved with $L=4$. When the SNR is smaller than $5$ dB, the accuracies are smaller than $60\%$ for all $L$. The accuracy increases rapidly when the SNR increases from 00 dB to $10$ dB and reaches an upper bound at about $20$ dB. This is because larger SNR results in lower RPCD reconstruction errors, which offers higher classification accuracy according to Fig. 4 and Fig. 8.
@@ VI-D Material Classification Accuracy versus Bandwidth
With SNR = $30$ dB and under the same other simulation settings as in subsection C, we explore the material classification accuracy versus the bandwidth by changing the number of subcarriers $K$ while keeping the frequency step $\Delta f$ unchanged, as shown in Fig. 9. It is seen that, the classification accuracy increases with the increase of $K$ in all cases. The highest accuracy is achieved at $L=4$ and $K=64$, where the accuracy is $88\%$. When $K$ is larger than $32$, the accuracies are larger than $50\%$ for all $L$. Generally speaking, larger $K$ leads to lower RPCD reconstruction errors, which further provides higher classification accuracy according to Fig. 6 and Fig. 9.
VII Conclusion

### v2

@@ 3. Closed and convex constraint set:
As shown in Fig. 2, the different subcarriers provide distinct pilot patterns in region $D$. However, due to the relatively small number of the UE antennas in practice, the stripes of the patterns generated by the same UE are roughly in the same direction that is mainly determined by the position of the UE. Thus, a single UE may not be sufficient to generate pilots for the reconstruction of RPCD within $D$. To address this problem, we require multiple users at different positions to provide diverse views and resolutions of RPCD. Besides, since UE-2 is closer to $D$, the near-field effect is more prominent and the pilot pattern stripes are not parallel. Closer proximity leads to more diversity in the pattern style, which is more beneficial to RPCD reconstruction. Despite the fact that each UE designs its pilots independently in (48), the pilot patterns of different UEs are inherently distinct and with low mutual coherence when the UEs are in different positions.
Remark 5.
OFDM contributes in reducing mutual coherence by enlarging the column space of $[\tilde{\mathbf{H}}_{1,1,u},\cdots,\tilde{\mathbf{H}}_{1,K,u}]$. No matter how large $I$ is, the dimension of the column space of $\tilde{\mathbf{H}}_{1,k,u}$ for each $k$ is bounded by $\text{dim}(\text{col}(\tilde{\mathbf{H}}_{1,k,u}))=\text{dim}(\text{col}(\mathbf{H}_{1,k,u}\tilde{\mathbf{W}}_{k,u}))\leq N_{t}$ due to $\mathbf{H}_{1,k,u}\in\mathbb{C}^{M\times N_{t}}$, which is usually much smaller than $M$. However, the dimension of the column space of the combined matrix $[\tilde{\mathbf{H}}_{1,1,u},\cdots,\tilde{\mathbf{H}}_{1,K,u}]$ with $K$ subcarriers is bounded by $\text{dim}(\text{col}([\tilde{\mathbf{H}}_{1,1,u},\cdots,\tilde{\mathbf{H}}_{1,K,u}]))\leq N_{t}K$. The dimension of the column space increases significantly from not larger than $N_{t}$ to potentially $N_{t}K$, which effectively enlarges the space of pilot patterns generated by the $u$-th UE. Consequently, the expansion of the column space reduces the likelihood of overlap between pilot patterns from the $u$-th UE, thereby decreasing mutual coherence in the OFDM sensing system.
@@ V-B Material Identification Methodology
In order to determine the material of the target, we first need to distinguish between the parts of region $D$ occupied by different materials. To accomplish this, we utilize the density-based spatial clustering of applications with noise (DBSCAN) algorithm to divide the sampling points within $D$ into several categories [52]. DBSCAN is an unsupervised algorithm with strong generalization of clusters for different shapes and sizes, and is guaranteed to converge regardless of the input data distribution. The approach to determining the DBSCAN hyper-parameters has been discussed in [53]. Since the relative permittivity and conductivity have different dimensions, we adopt the non-dimensional and scale-invariant Mahalanobis distance as the metric in the DBSCAN clustering algorithm [54]. The number of clusters is self-adaptive and represents the estimated number of material categories. The cluster centroid of the air, which represents the average permittivity and conductivity values of the air, is typically close to $(1,0)$. Conversely, the cluster centroids of the target materials, which individually represent their average permittivity and conductivity, are generally positioned significantly away from $(1,0)$.
@@ VI Simulation Results and Analysis
Suppose the target is located within the region $D=[-1,1]\times[-1,1]$ $\mathrm{m}^{2}$ unless otherwise specified. A total of $U=10$ UEs are randomly located in a circular region of radius $10$ m centered at the origin but outside region $D$. We consider a ULA transmitter equipped with $N_{t}=8$ antennas at each UE. The power budget is the same for all UEs and is evenly allocated to all subcarriers. For each UE, $I=16$ pilot OFDM symbols are transmitted on each subcarrier. Assume $4$ BSs are located at $(100,0)$ m, $(0,100)$ m, $(-100,0)$ m, and $(0,-100)$ m for $l=1,2,3,4$, respectively, and are fused in sequence. We set $\zeta_{l}=1/L,l=1,\cdots,L.$ Each BS is equipped with $N_{r}=64$ ULA antennas, and the ULA is perpendicular to the line from the origin to the BS center. We assume that the sum of thermal noise and possible interference from unknown EM sources or scatters at each BS is subject to the same zero-mean circularly symmetric complex Gaussian distribution with the same covariance. SNR is then defined as the ratio of the overall received signal power from all BSs to the overall noise power. The central frequency is $f_{c}=28$ GHz and the frequency spacing of OFDM subcarriers is $\Delta f=240$ KHz according to the 3GPP standard [57]. The inter-antenna spacing for both the UEs and the BSs is set as $\lambda_{c}/2=0.0054$ m. We choose the number of sampling points in region $D$ as $M=64\times 64=4096$. Each sampling point within $D$ can be regarded as a single pixel in an image, i.e., the RPCD image consists of $64\times 64=4096$ pixels.
@@
(b) Reconstructed relative permittivity
@@
(c) Reconstructed relative permittivity
@@
(e) Reconstructed conductivity with SNR = 0 dB
(f) Reconstructed conductivity with SNR = 30 dB
Figure 3: RPCD reconstruction results versus SNR with $L=4$ BSs and $K=64$ subcarriers. Unit of conductivity is S/m.
Figure 4: NMSE of PRCD versus SNR with $K=32$.
@@ VI-A RPCD Reconstruction Performance versus SNR
To illustrate the RPCD reconstruction results vividly, we present the reconstructed images of the target based on relative permittivity and conductivity, respectively, with $L=4$, $K=32$, and SNR $=0$ dB or $30$ dB in Fig. 3. The target “THU” is the abbreviation of “Tsinghua University”, and the three letters are composed of wood, chipboard, and plasterboard, respectively. It is seen from Fig. 3 that, the reconstructed RPCD can reflect the general shape of the target. The RPCD reconstructed at SNR $=30$ dB is much more accurate to demonstrate the target’s shape compared to the RPCD reconstructed at SNR $=0$ dB. Moreover, a higher SNR value results in more accurate reconstructed values of relative permittivity and conductivity, which leads to better representation of the target’s real EM property.
We explore the NMSE of reconstructed RPCD versus the SNR, as shown in Fig. 4. We set the number of subcarriers to be $K=32$ and sequentially select several BSs and fuse their data at the feature level to enhance the overall accuracy of the reconstructed RPCD. It is seen from Fig. 4 that, the NMSE decreases with the increase of SNR in all cases, and the best performance is achieved when $L=4$. The NMSE decreases fast in the beginning when the SNR increases from 00 dB to $10$ dB. When the SNR reaches a threshold of approximately $20$ dB, the NMSE reduces to an error floor and remains largely unchanged. At this level, the limitation on RPCD reconstruction quality is no longer due to noise at the BS but rather to the insufficient information about the target captured in the general sensing matrix. As signals from more BSs are utilized, more information of the target is integrated, which leads to a smaller error floor in RPCD reconstruction.
@@ VI-B RPCD Reconstruction Performance versus Bandwidth
We investigate the NMSE of reconstructed RPCD versus the bandwidth by changing the number of subcarriers $K$ while keeping the frequency step $\Delta f$ unchanged, as shown in Fig. 6. It is seen that, NMSE decreases with the increase of $K$ in all cases, and the best performance is achieved when the signals from all 4 BSs are used. This is because the increased bandwidth enhances the diversity of the pilot patterns and reduces the mutual coherence of the sensing matrix. Consequently, more information about RPCD is extracted, leading to better reconstruction accuracy.
@@ VI-C Material Classification Accuracy versus SNR
Figure 8: Material classification accuracy versus SNR with $K=32$.
@@
To provide an example of classifying the materials in region $D$, we show the clustering results using DBSCAN in Fig. 7 with $K=32$, SNR = $30$ dB, and $L=4$. In Fig. 7, each data point represents a sampling point in region $D$, and the colors of the data points represent the classification results. It is observed that the cluster of the air is close to $(1,0)$, whereas the clusters of the target materials are notably away from $(1,0)$. Therefore, this separation enables clear differentiation between sampling points associated with air and those associated with the target materials.
Using the sampling points occupied by the target after removing the outlier points and the points associated with air, we show the material classification accuracy versus SNR with different $L$ and $K=32$ in Fig. 8. It is seen that, the classification accuracy increases with the increase of SNR for all $L$, and the best performance is achieved with $L=4$. When the SNR is larger than $15$ dB, the accuracies are larger than $45\%$ for all $L$. The accuracy increases rapidly when the SNR increases from 00 dB to $10$ dB and reaches an upper bound at about $20$ dB. This is because larger SNR results in lower RPCD reconstruction errors, which offers higher classification accuracy according to Fig. 4 and Fig. 8.
@@ VI-D Material Classification Accuracy versus Bandwidth
With SNR = $30$ dB and under the same other simulation settings as in Section V.C, we explore the material classification accuracy versus the bandwidth by changing the number of subcarriers $K$ while keeping the frequency step $\Delta f$ unchanged, as shown in Fig. 9. It is seen that, the classification accuracy increases with the increase of $K$ in all cases. When $K$ is larger than $35$, the accuracies are larger than $50\%$ for all $L$. Generally speaking, larger $K$ leads to lower RPCD reconstruction errors, which further provides higher classification accuracy according to Fig. 6 and Fig. 9.
VI-E Influence of Channel Error
Figure 10: NMSE of PRCD versus NMSE of $\mathbf{H}_{1,k,u}$ with SNR = $30$ dB and $K=32$.
Figure 11: NMSE of PRCD versus NMSE of $\mathbf{H}_{2,k,l}$ with SNR = $30$ dB and $K=32$.
We explore the influence of the channel error by adding CSCG noise to the computed channels $\mathbf{H}_{1,k,u}$ and $\mathbf{H}_{2,k,l}$. The NMSE of PRCD versus NMSE of $\mathbf{H}_{1,k,u}$ with SNR = $30$ dB and $K=32$ is shown in Fig. 10. It is seen that as the NMSE of $\mathbf{H}_{1,k,u}$ increases, the NMSE of the RPCD also rises for all values of $L$. Systems with more BSs consistently demonstrate better performance, exhibiting lower NMSE values of PRCD. An error floor appears when the NMSE of $\mathbf{H}_{1,k,u}$ is between -40 dB and -25 dB, where the NMSE curves flatten out for all values of $L$. In this range, reducing the channel error further does not significantly improve the NMSE of PRCD. While the proposed method shows robustness to small channel errors of $\mathbf{H}_{1,k,u}$, increasing the number of BSs does not necessarily enhance its resilience to larger channel errors.
We explore the NMSE of PRCD versus NMSE of $\mathbf{H}_{2,k,l}$ with SNR = $30$ dB and $K=32$ in Fig. 11. We observe phenomena similar to Fig. 10, but with smaller NMSE values under the same NMSE of channel error. This improvement may be attributed to the fact that the BSs are relatively farther from the target than the UEs, which causes the channel $\mathbf{H}_{2,k,l}$ to approximate a far-field single-path model. As a result, the effective degrees of freedom for the channel $\mathbf{H}_{2,k,l}$ are smaller than $\mathbf{H}_{1,k,u}$, which makes the sensing system more resilient to the errors of $\mathbf{H}_{2,k,l}$. The proposed method demonstrates robustness to small channel errors of $\mathbf{H}_{2,k,l}$ with varying numbers of BSs.
VI-F Influence of Target’s Location
Figure 12: NMSE of PRCD versus $K$ with SNR = $30$ dB and $D=[49,51]\\ \times[49,51]$ $\mathrm{m}^{2}$.
We explore the influence of the target’s location by changing the region of the target into $D=[49,51]\times[49,51]$ $\mathrm{m}^{2}$. The $10$ UEs are randomly located in a circular region of radius $10$ m centered at $[50,50]$ m but outside region $D$. The NMSE of PRCD versus $K$ with SNR = $30$ dB is shown in Fig. 12, where the $4$ BSs located at $(100,0)$ m, $(0,100)$ m, $(-100,0)$ m, and $(0,-100)$ m are fused in sequence for $L=1,2,3,4$, respectively. We observe phenomena similar to Fig. 6 where the target region is at the center of multiple BSs, but with relatively smaller NMSE values under the same $K$. The reason is that the BSs at $(100,0)$ m and $(0,100)$ m are closer to the target when the target is located within $D=[49,51]\times[49,51]$ $\mathrm{m}^{2}$ than when the target is located within $D=[-1,1]\times[-1,1]$ $\mathrm{m}^{2}$. The BSs at $(100,0)$ m and $(0,100)$ m can thus provide richer information of the target’s EM property, which contributes to smaller NMSE values. Moreover, when the target is located within $D=[49,51]\times[49,51]$ $\mathrm{m}^{2}$, the BSs at $(100,0)$ m and $(0,100)$ m are closer to the target and thus have smaller NMSE of PRCD reconstruction than the BSs at $(-100,0)$ m and $(0,-100)$ m. Adding the information collected by the BSs at $(-100,0)$ m and $(0,-100)$ m to the information collected by the BSs at $(100,0)$ m and $(0,100)$ m results in relatively small improvement in PRCD reconstruction. Thus, the curves corresponding to $L=2,3,4$ are closer to each other with $D=[49,51]\times[49,51]$ $\mathrm{m}^{2}$ than the curves corresponding to $L=2,3,4$ with $D=[-1,1]\times[-1,1]$ $\mathrm{m}^{2}$ in Fig. 6.
VII Limitations, Potential Challenges, Practical Implementation, and Future Directions
The proposed EM property sensing framework, while effective, has several limitations. The framework assumes static environments and is designed to sense static target. Moreover, the multi-BS data fusion adds extra computational overhead, which may present challenges in resource-constrained environments.
Potential challenges include handling high-dimensional data and ensuring real-time processing, especially as more BSs are added. Accurate sensing and synchronization across multiple BSs also pose potential challenges, especially in dynamic environments with moving targets.
The practical implementation of the proposed EM property sensing framework requires both advanced hardware and software components. On the hardware side, multiple BSs must be equipped with multi-antenna arrays and have robust communication links. Additionally, high-performance processors are necessary to handle the computational complexity of data fusion and signal processing, especially for real-time applications. On the software side, the system requires efficient algorithms related to pilot design, compressive sensing, and data fusion for EM property sensing and material classification. Real-time operating systems or parallel computing frameworks would be essential to meet the processing speed and resource management demands.
Future directions include extending the framework to sense the EM property of moving targets, which could be beneficial in areas like surveillance and healthcare. Incorporating machine learning techniques for enhanced material identification and exploring adaptive sensing strategies to handle dynamic conditions are promising avenues. Moreover, improving robustness to noise and CSI errors, along with refining multi-BS fusion techniques, could significantly enhance the framework’s real-world applicability. Potential improvements in pilot design could involve designing pilots that simultaneously perform both channel estimation and EM property sensing, optimizing resource use. Fusion algorithms could be enhanced by dynamically adjusting the fusion weights across BSs, further improving performance. Additionally, the proposed MACE method could be extended to other sensing tasks that can be formulated as the optimization of a sum of functions constructed by BSs, such as localization or velocity estimation, making it versatile for a wide range of applications.
VIII Conclusion